# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v3

@@ Abstract
The accomplished functions of a variety of quantum devices are closely associated with the controlling of exciton transport. To this end we study the exciton transport of the two-dimensional system consisting of two-level multichains with various coupling configurations in a cavity. Two types of the chains, the homogeneous and heterogeneous coupling chain, as well as two inter-chain coupling conformations, the square and triangle type, are considered. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on exciton transport are in detail investigated for different coupling configurations through the spectra, the Hopfield coefficients, and the steady-state dynamics of the system. The results show that in the absence of cavity the exciton transport currents and efficiency are decided by the distribution of exciton on the multichains. However, when the cavity is considered the exciton transport can be significantly enhanced or suppressed by the polariton at the cavity-dressed energy level crossings and anticrossings near zero-energy modes, at which the coherent excitation and Landau-Zener transitions occur. Meanwhile, we discover that the incontinuity and extreme points of the second order partial derivative of the Hopfield coefficients of the photon with the inter-chain coupling and the dimerization parameter correspond respectively to the crossings and anticrossings at the extreme points of the photon occupation number. In addition, the exciton transport efficiency is closely related to the parity of both the length and number of the chains, which behaves as oscillation. This work is important for understanding the exciton transport mechanism in the multichain-cavity system, and provides theoretical basis for excitonic devices with controllable and efficient exciton transport.
@@ I Introduction
In quantum device, the carriers of energy and information transport are mainly electron, photon, and exciton. Enhancing and controlling the efficiency, the speed, the distance, and the current of carrier transport are the center to accomplish the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient ; hagenmuller2017cavity , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. The factors affecting exciton transport are various, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity ; allard2022disorder ; balasubrahmaniyam2023enhanced ; ribeiro2022multimode , defect sites wagner2011bound ; akselrod2014visualization , and topological property buendia2024long ; allard2023multiple ; nie2021topology . Thus, a large of studies are devoted to revealing the exciton transport mechanisms and properties of various quantum system engel2007evidence ; menke2013tailored , and to achieving the requirements for the property of exciton transport through reduce and avoid detrimental factors.
@@
In experiment, the one-dimensional (1D) system carrying electron, photon, or exciton transport can be realized by various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room and the polaritons forming between the edge state excitons and the cavity photons kavokin2003cavity ; agranovich2003cavity , in which the molecular energy structures are modified galego2015cavity . The presence of polaritons can adjust the efficiency wei2019enhanced ; wei2022cavity ; feist2015extraordinary , rate wang2021polariton ; aroeira2023theoretical , and distance du2018theory ; myers2018polariton ; sandik2024cavity ; takazawa2010fraction of the exciton transport because the decay of excitons on chains is suppressed.
@@
The hybrid system consisting of an 1D molecule chain and a cavity can be described by the Tavis-Cummings model tavis1968exact . The chain with the staggered hopping interaction is depicted by the Su-Schrieffer-Heeger (SSH) model su1979solitons . In the SSH chain there are topologically trivial and nontrivial phases for the even number of the two-level systems (TLSs) cai2019observation . The even chain in the topologically nontrivial regime has a small gap bewteen two localized nearly-zero-energy modes in the middle of the topological bandgap, which leads to the formation of the hybrid edge modes with the left and the right edge state. Therefore, the nontrivial edge states topologically protected can enhance the exciton transport for a finite-size SSH chain, which was verified theoretically and experimentally chang2020probing ; wei2022cavity . In fact, the nature of the gap is the energy level anticrossing, at which the Landau–Zener (LZ) transition occurs. In the two-dimensional (2D) and three-dimensional tungsten disulfide material, the anti-crossing dispersions were presented, in which the boosted transport capability of excitons was corroborated in experiment guo2022boosting ; yuan2017exciton . The properties of the multichains system coupling to a cavity were theoretically investigated kvande2023finite . Thus, the potential properties of the more than 1D system on the exciton transport are fascinating and valuable to study.
@@
Motivated by the works on the high dimensional tungsten disulfide, in this paper we devote to the studies of the inter-chain joint effect on exciton transport in 2D TLSs array with a cavity. Each chain contains two alternative coupling types, i.e., the homogeneous (HO) and the heterogeneous (HE) coupling, and two conformations of the inter-chain coupling are set, i.e., square and triangle types. Firstly, the energy spectra of the hybrid system in the condition of the strong exciton-cavity coupling with the inter-chain coupling strength and the dimerization parameter are presented by diagonalizing numerically the Hamiltonian for six configurations. Meanwhile, we calculate the von Neumann entropy of exciton to confirm the locations of the crossings and anticrossings, and the second order partial derivative of the photonic Hopfield coefficients to show the change of the occupation number of the cavity photon. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport are investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which its current and efficiency depend on the distribution of exciton on the whole multichains. Importantly, for the hybrid system the exciton transport is obviously related to the crossings and anticrossings of the dressed nearly-zero-energy modes due to the coherent excitation and LZ transition between the multichains and the cavity, at which the excitation is exchanged. As a consequence the exciton transport can be controlled by the crossings and anticrossings, which depend on the configurations of the multichains. Here the polaritons act as important role. In addition, the second order partial derivative of the Hopfield coefficients of the photon is a good physical quantity to measure the extreme points of the photon occupation number and the locations of the crossings and anticrossings. The optimal parameters regions on the spaces of the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter are presented. Further, the results show that the exciton transport dynamics depend on the odevity of the number and length of the chains.
@@
The remainder of this paper is organised as follows: In Sec. II the theoretical model is introduced including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), the formula calculating the von Neumann entropy of excitons (in Sec. II.2), and the Lindblad quantum master equation of the system with external driving field (in Sec. II.3). In Sec. II.4 the exciton transport currents and efficiency for the steady state are derived. In Sec. III the energy spectrums are given by numerically diagonalizing the Hamiltonian. In Sec. IV we present and discuss the steady-state dynamics of excitons for various configurations of the multichains, in which the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport are analysed. Section V is a brief summary and future perspective.
@@
Figure 1: (color online) The schematic representation of the hybrid system that the 2D TLSs array are placed in a cavity for the three configurations HO ST (a), HO TT (b), and HE ST (c). The red sphere represents the TLS with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and via the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. The intra-chain coupling strength is $h$. $\delta$ is the dimerization strength of HE chain. It exists the inter-chains couplings with the strength $\zeta$ between the nearest TLSs in two nearest chains.
@@ II.1 Hamiltonian
The considered system is a 2D molecular aggregate placed in a single-mode cavity, which can be modelled as multichains with $L$ ($L\geqslant 2$) chains as shown in Fig. 1. The $k$th chain has $N_{k}$ identical TLSs. For simplicity, we set that the numbers of the TLSs of each chain are the same ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupling system can be expressed as
@@
where $H_{{\rm e},k}$ are the Hamiltonian of the $k$th bare chain.
Considering the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the SSH model as
@@
where $h$ characterizes the hopping amplitude between two nearest-neighbor TLS monomers in a chain, $\delta$ represents the dimerization strength, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ (setting $\hbar=1$) at the $i$th site of the $k$th chain with the groud state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$ because of the identical partical, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, ${\rm H.c.}$ denotes the Hermitian conjugate. For $\delta=0$ and $\delta\neq 0$, the SSH chain behaves entirely different property, which correspond respectively homogeneous and heterogeneous coupling in a molecular chain. In addition, it exist the topologically nontrivial regime for $\delta\neq 0$. In order to emphasize the difference of the exciton transport of the two coupling cases, we designated separately them as the homogeneous (HO) and the heterogeneous (HE) coupling chain in the next content.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest types of the configuration, which are named as the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
@@
and $\zeta$ denotes the inter-chain coupling strength, namely the interaction strength between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multichains have six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as HO ST, HO TT, HE ST, HE TT, HOCC-HE ST, and HO-HE TT. The HO ST, HO TT, and HE ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. Correspondingly, the HE TT is that the HO chains in the HO TT are substituted by the HE chains. Particularly, the HO and HE chain are arranged with staggered form in HO-HE ST and HO-HE TT, in which we set that the odd chains are HO and the even chains are HE.
In the rotating frame of the driving frequency $\omega$, the radiation part of cavity is described by $H_{\rm c}$ as
@@
where $c^{\dagger}=|c\rangle\langle 0|$ ($c=|0\rangle\langle c|$) is the photon creation (annihilation) operator, the cavity mode is decided by the frequency $\omega_{\rm c}$, $\Delta_{\rm c}=\omega-\omega_{\rm c}$ is the detuning of the cavity field frequency $\omega_{\rm c}$ with respect to the $\omega$, $|c\rangle$ is the single-photon state that only the cavity field is excited, and $|0\rangle$ is the vacuum state of the exciton-cavity system. In order to satisfy the resonant condition between the cavity and the TLSs, we set $\omega_{\rm c}=\omega_{\rm a}$, hence $\Delta_{\rm a}=\Delta_{\rm c}=\Delta$.
In the rotating frame of $\omega$ the coupling between the multi-chain aggregate and the cavity is given by
@@
where $g_{k,i}$ describes the coupling strength between the TLS at the $i$th site in the $k$th chain and the cavity field. For the sake of simplicity, we assume that $g_{k,i}$ is uniform for all TLSs, that is $g_{k,i}=g$. Generally, $g$ depends on the collective Rabi splitting $\Omega_{\rm R}=2\sqrt{\sum_{k=1}^{L}\sum_{i=1}^{N}g^{2}_{k,i}}=2g\sqrt{LN}$.
In order to investigate the excitation dynamics, we introduce a external coherent driving in the exciton-cavity system. In this paper we choose two driving means, i.e., the driving field is exerted on the first site of the $k$th chain as well as simultaneously on this site and the cavity, which can be described respectively as
@@
where $\omega$ is the driving frequency, $\xi_{{\rm T},k}$ ($\xi_{\rm c}$) is the strength of the TLS (cavity) drive. The second mean is denoted by ”TLS+cavity” in the following. Here, we set $k=1$, namely, only the first site in the first chain is driven. The two driving forms were used to study the two-photon blockade in the cavity QED system hamsen2017two . In the rotating frame of $\omega$, the Eq. (9) can be written as
@@
For the hybrid system consisting of excitons and photons, its eigenstates can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to the conclusion of Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, at which the LZ transitions occur. The limiting of distance between two anti-cross energy levels is crossing that means the energy degeneracy. At the crossings, the excitations are exchanged coherently between the multichains and the cavity. As a result, in energetically degenerate state the exciton transport can be observed tagarelli2023electrical and boosted lin2012efficient .

### v4

@@ Abstract
The performance of various quantum devices is fundamentally linked to the control of exciton transport. To explore this, we study the exciton transport of the two-dimensional multi-chain systems with different coupling configurations in an optical cavity. Two types of the chains—the homogeneous and heterogeneous coupling chain, as well as two inter-chain coupling conformations—the square and triangular arrangements, are considered. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on exciton transport are systematically investigated for different coupling configurations through the spectra, the Hopfield coefficients, and the steady-state dynamics of the system. The results show that in the absence of a cavity the exciton transport currents and efficiency are determined by the exciton distribution across the multi-chain system. However, when a cavity is introduced the exciton transport can be significantly enhanced or suppressed by the polariton formation at the cavity-dressed energy level crossings and anticrossings near zero-energy modes, where the coherent excitation and Landau-Zener transitions occur. Meanwhile, we discover that the discontinuous and extremal points in the second-order partial derivatives of the photon Hopfield coefficients with respect to the inter-chain coupling and the dimerization parameter correspond respectively to the crossings and anticrossings at the extreme points of the photon occupation number. Additionally, exciton transport efficiency is closely related to the odevity of both chain length and chain number, and exhibits oscillatory behaviour. This work provides critical insights into the exciton transport mechanism in multi-chain–cavity system and theoretical basis for designing high-performance excitonic devices with tunable transport properties.
@@ I Introduction
In quantum device, energy and information are primarily transported by three carriers—electron, photon, and exciton. Optimizing and controlling the efficiency, the speed, the distance, and the current of carrier transport are the center to realize the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient ; hagenmuller2017cavity , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. Exciton transport is influenced by multiple factors, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity ; allard2022disorder ; balasubrahmaniyam2023enhanced ; ribeiro2022multimode , defect sites wagner2011bound ; akselrod2014visualization , and topological property buendia2024long ; allard2023multiple ; nie2021topology . Consequently, extensive research focuses on revealing the exciton transport mechanisms and properties in diverse quantum systems engel2007evidence ; menke2013tailored , and improving transport properties by reducing and avoiding detrimental factors.
@@
In experiment, the one-dimensional (1D) system carrying electron, photon, or exciton transport have been realized through various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room , in which the polaritons form kavokin2003cavity ; agranovich2003cavity , and the molecular energy structures are modified galego2015cavity . The polariton formation significantly modulates exciton transport properties containing the efficiency wei2019enhanced ; wei2022cavity ; feist2015extraordinary , rate wang2021polariton ; aroeira2023theoretical , and distance du2018theory ; myers2018polariton ; sandik2024cavity ; takazawa2010fraction because the decay of excitons on chains is suppressed.
@@
The hybrid system composed of a 1D molecular chain coupled to an optical cavity is well-described by the Tavis-Cummings model tavis1968exact . The chain with the staggered hopping interaction is depicted by the Su-Schrieffer-Heeger (SSH) model su1979solitons , which exhibits topologically trivial and nontrivial phases for even-numbered the two-level systems (TLSs) cai2019observation . In the topologically nontrivial phase, even-length chains host a small gap bewteen two localized nearly-zero-energy modes within the topological bandgap, enabling hybrid edge modes formation from the left and the right edge states. These topologically protected edge states enhance exciton transport in finite SSH chains, as confirmed theoretically and experimentally chang2020probing ; wei2022cavity . The gap’s origin lies in energy-level anticrossings, where the Landau–Zener (LZ) transitions occur. In the two-dimensional (2D) and three-dimensional tungsten disulfide material, the anti-crossing dispersions were presented, in which the enhancement of exciton transport capability was corroborated in experiment guo2022boosting ; yuan2017exciton . Theoretical study have explored cavity-coupled system with multichains kvande2023finite , suggesting that the exciton transport properties of the higher-dimensional ($\geq 1$) systems warrant further investigation.
@@
Motivated by the recent studies on the high dimensional tungsten disulfide, in this paper we devote to the studies of the inter-chain joint effect on exciton transport in 2D TLSs array within a cavity. Each chain contains two alternative coupling types, i.e., the homogeneous (HO) and the heterogeneous (HE) coupling, and two conformations of the inter-chain coupling are set, i.e., square and triangle types. Firstly, the energy spectra of the strong exciton-cavity coupled system are obtained by diagonalizing numerically the Hamiltonian for six configurations, analyzing their dependence on inter-chain coupling strength and dimerization parameter. Meanwhile, we calculate the von Neumann entropy of exciton to confirm the locations of the crossings and anticrossings, and the second order partial derivative of the photon Hopfield coefficients to show the change of the occupation number of the cavity photon. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport are investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which its current and efficiency depend on the distribution of exciton on the whole multichains. Importantly, the exciton transport in the hybrid system correlates strongly with the crossings/anticrossings of dressed near-zero-energy mode, where the coherent excitation and LZ transition mediate the excitation exchange between the multichains and the cavity. Notably, polaritons play pivotal roles in controlling exciton transport through the crossings/anticrossings. In addition, the second order partial derivatives of the photon Hopfield coefficients serve as reliable metrics for identifying the extremal points in photon occupation number and the precise locations of the crossings/anticrossings. Further, we present the optimal parameters regions with the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter. Finally, the exciton transport dynamics exhibit the odevity-dependent oscillations tied to the number and length of the chains.
@@
The remainder of this paper is organised as follows: Sec. II presents the theoretical framework, including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), Von Neumann entropy formulation for excitons (in Sec. II.2), the Lindblad quantum master equation of the system with an external driving field (in Sec. II.3), and the steady-state exciton transport currents and efficiency (in Sec. II.4). In Sec. III the energy spectra are given by numerically diagonalizing the Hamiltonian. In Sec. IV we simulate the steady-state exciton dynamics across multi-chain configurations, analyzing the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport. Section V is a brief summary and future perspective.
@@
Figure 1: (color online) The schematic representation of the hybrid system consisting of the 2D TLSs array within a cavity for three configurations HO ST (a), HO TT (b), and HE ST (c). Red spheres represent the TLSs with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and via the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. $h$ is the intra-chain hopping strength between adjacent TLSs. $\delta$ is the dimerization strength of HE chain. $\zeta$ is the inter-chains couplings strength between the nearest TLSs.
@@ II.1 Hamiltonian
The considered system is a 2D molecular aggregate embedded in a single-mode cavity, which can be modelled as multiple parallel chains with number $L$ ($L\geqslant 2$) as illustrated in Fig. 1. The $k$th chain comprises $N_{k}$ identical TLSs. For simplicity, we assume that all chains contain the same number of the TLSs ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupled system can be expressed as
@@
where $H_{{\rm e},k}$ are the term of the $k$th bare chain.
Considering the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the SSH model
@@
where $h$ characterizes the hopping amplitude between two nearest-neighbor TLS monomers in a chain, $\delta$ represents the dimerization strength, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ (setting $\hbar=1$) at the $i$th site of the $k$th chain with the ground state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$ because of the identical particle, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, and ${\rm H.c.}$ denotes the Hermitian conjugate. The SSH chain behaves entirely different property for $\delta=0$ and $\delta\neq 0$ corresponding to homogeneous and heterogeneous coupling in a molecular chain respectively, because the topologically nontrivial regime exists for the latter. In order to emphasize the difference in exciton transport between the two coupling cases, we designate separately them as the homogeneous (HO) and the heterogeneous (HE) coupling chain in the following content.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest configurations, the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
@@
and $\zeta$ denotes the inter-chain coupling strength, namely the hopping amplitude between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multi-chain system exhibits six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as HO ST, HO TT, HE ST, HE TT, HOCC-HE ST, and HO-HE TT. The HO ST, HO TT, and HE ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. The HE TT configuration is obtained by replacing the HO chains in HO-TT with HE chains. For the hybrid cases, the HO and HE chains are arranged in a staggered pattern in both HO-HE ST and HO-HE TT, in which the odd-numbered and even-numbered chains are respectively HO and HE type.
In the rotating frame of the driving frequency $\omega$, the cavity radiation field is described by the Hamiltonian
@@
where $c^{\dagger}=|c\rangle\langle 0|$ ($c=|0\rangle\langle c|$) represents the photon creation (annihilation) operator, $\omega_{\rm c}$ is the cavity-mode frequency, $\Delta_{\rm c}=\omega-\omega_{\rm c}$ is the detuning of $\omega_{\rm c}$ with respect to $\omega$, $|c\rangle$ denotes the single-photon excited state of the cavity field, and $|0\rangle$ represents the joint vacuum state of the exciton-cavity system. To ensure the resonant condition between the cavity and the TLSs, we set $\omega_{\rm c}=\omega_{\rm a}$, hence $\Delta_{\rm a}=\Delta_{\rm c}=\Delta$.
In the rotating frame of $\omega$, the light-matter coupling between the multi-chain aggregate and the cavity is given by
@@
where $g_{k,i}$ represents the coupling strength between the TLS at the $i$th site in the $k$th chain and the cavity field. For the sake of simplicity, we assume uniform coupling $g_{k,i}=g$ for all TLSs. Generally, $g$ depends on the collective Rabi splitting $\Omega_{\rm R}=2\sqrt{\sum_{k=1}^{L}\sum_{i=1}^{N}g^{2}_{k,i}}=2g\sqrt{LN}$.
In order to investigate the excitation dynamics, we introduce an external coherent drive in the exciton-cavity system. We consider two driving means, i.e., the driving field is exerted on the first site of the $k$th chain, as well as simultaneously on both this site and cavity, which can be described respectively by
@@
where $\omega$ is the driving frequency, $\xi_{{\rm T},k}$ ($\xi_{\rm c}$) denotes the driving strength for the TLS (cavity). The second mean is denoted by ”TLS+cavity” in the following. Here, we set $k=1$, namely, only the first site in the first chain is driven. The two driving means were used to study the two-photon blockade in the cavity QED system hamsen2017two . In the rotating frame of $\omega$, the Eq. (9) can be written as
@@
For the exciton-photon hybrid system, its eigenstates can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maxima or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, where the LZ transitions occur. The transition from anticrossing to crossing means the energy degeneracy hu2018effect . At the crossings, the excitations are exchanged coherently between multichains and cavity. As a result, the boosted exciton transport in degenerate states was demonstrated tagarelli2023electrical ; lin2012efficient .